OpenStudy (anonymous):

Find the value of y for which the line through A and B has the given slope m: A(-2, 3), B(4, y), m=-2/3.

OpenStudy (anonymous):

I know, but my stupid professor give us assignments like this. And I can't memorize most of these stuffs.

OpenStudy (anonymous):

When did I laugh at your question?

OpenStudy (anonymous):

Whatever, you win. Do you know how to do this problem?

OpenStudy (anonymous):

I got -1 but that's not the right answer.

OpenStudy (anonymous):

You're supposed to solve for y, the slope is already given, -2/3. (y-3)/(4+2)=-2/3. And I got y=-1. Is this right?

OpenStudy (anonymous):

y in B(4, y).

OpenStudy (anonymous):

Then how come the answer key says something like y=-5/3?

OpenStudy (anonymous):

That stupid professor must die from my hand someday, I'm serious. Anyways, thanks, mean woman.

OpenStudy (anonymous):

Who's stubborn?

OpenStudy (anonymous):

HAHAHA, you're not stubborn. I'm stubborn.

OpenStudy (anonymous):

OK. hehehe.

OpenStudy (loser66):

@Microrobot

OpenStudy (anonymous):

Idealist can you provide some more information

OpenStudy (anonymous):

What is it you have done to try to get the answer

OpenStudy (loser66):

oops, sorry friend, she is offline.

OpenStudy (anonymous):

Oh, it's ok

OpenStudy (anonymous):

Wait 5 minutes.

OpenStudy (loser66):

@Idealist he is here to help us and the person who need help are us. We, the askers, cannot ask the helper to wait for us, friend.

OpenStudy (anonymous):

I'm sorry. You're right, it's my fault. Next time I wouldn't be like this. I had special things to do, that's why.

OpenStudy (loser66):

now, his turn, he is offline, we cannot get his help, let me tag others

OpenStudy (loser66):

@jim_thompson5910

OpenStudy (anonymous):

It's alright. Please don't do it anymore. I'll get help from others tomorrow.

OpenStudy (loser66):

ok, it's ok to me.